OpenStudy (anonymous):

Points (-3, y) and (5, -6) lie on a line with slope -5/4. Find the value y.

OpenStudy (anonymous):

you have the equation for the slope \[m = \frac{ y _{1} - y _{2} }{ x_{1} - x_{2} }\] from this, you can extract y2: \[y_{1} = y_{2} + (x_{1} - x_{2})*m\] plug in the numbers and you'll get \[y = -6 + (-3 - 5)*\frac{ -5 }{ 4 } = 4\]

OpenStudy (anonymous):

Thank you(: